OpenStudy (anonymous):

find the derivative of the function: y = 1/(5x-7)^7

OpenStudy (jdoe0001):

\(\bf y=\cfrac{1}{(5x-7)^7}\implies y=(5x-7)^{-7}\) chain rule it

OpenStudy (anonymous):

but dont u need to use the quotient rule also?

OpenStudy (jdoe0001):

sure, you could use the quotient rule as well longer way, but yes

OpenStudy (anonymous):

how did the power 7 became negative?

OpenStudy (akashdeepdeb):

\[y = \frac{1}{(5x-7)^7}\] \[y = (5x-7)^{-7}\] Use chain rule now. \[\frac{dy}{dx} = \frac{d((5x-7)^{-7})}{dx} = -7(5x-7)^{-8}.5 = -35(5x-7)\]

OpenStudy (jdoe0001):

\(\Large a^{-{\color{red} n}} = \cfrac{1}{a^{\color{red} n}}\)

OpenStudy (anonymous):

\[-35(5x - 7)^{-8}\] would that be the answer?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

thank you, god bless

OpenStudy (jdoe0001):

yw